# Quantifying uncertainty in climate projections with conformal ensembles

### v2

@@ 3.3 Conformal prediction sets for climate fields
Given a sufficiently adequate multi-model ensemble analysis function $f_{\theta}\in\mathcal{F}_{\Theta}$, where $f_{\theta}:\mathbbm{R}^{p}\mapsto\mathbbm{R}^{q}$, and a valid depth function $d:\mathbbm{R}^{q}\times\mathcal{P}\mapsto[0,1]$ (Equation 4), we can implement the general ICP algorithm (Algorithm 1) for climate projection. Using the notation from Section 3.1, suppose we have a historical dataset $D_{\text{hist}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, an analysis function $f_{\theta}$, a depth function $d$, a confidence level $\alpha\in(0,1)$, and an out of sample input $X_{n+1}$. Algorithm 2 describes our procedure to compute $C_{\alpha}(X_{n+1})$ such that $P(Y_{n+1}\in C_{\alpha}(X_{n+1}))\geq 1-\alpha$.
@@
Algorithm 2 can be interpreted as “converting” the multi-model ensemble of climate model runs observed at time $n+1$, $X_{n+1}$, into an exact conformal prediction set $C_{\alpha}(X_{n+1})$. We take an inexact approximation of the projection distribution, model ensembles, and convert it to a more rigorous approximation, conformal ensembles. This approach is similar to Conformalized Quantile Regression (CQR) methods (Romano, Patterson and Candes, 2019; Kivaranovic, Johnson and Leeb, 2019) that convert arbitrary prediction quantiles into exact conformal quantiles. However, because the depth function $d(\cdot)$ measures outlyingness over the entire spatial domain simultaneously, the resulting projection sets $C_{\alpha}(X_{n+1})$ are jointly valid over $Y_{n+1}$ instead of only pointwise valid.
@@ 3.4 Conformal ensembles for uncertainty quantification
The standard approach to representing multivariate or functional prediction sets $C_{\alpha}(X_{n+1})$, generated by Algorithm 2, is to compute the pointwise minimum and maximum bounds that form a prediction band (Lei, Rinaldo and Wasserman, 2015; Diquigiovanni, Fontana and Vantini, 2021). Bands are commonly used because they are computationally efficient and form a convex hull around the level sets of $\mathcal{R}_{cal}$. However, because bands are a loose approximation, they contain many projections that do not belong to $C_{\alpha}(X_{n+1})$, and they discard information about the overall projection density. This makes bands less than ideal for climate projection, where we do not want our prediction sets to contain highly implausible climate fields and we want to assess the projection density.
@@
$E_{\alpha}(X_{t})=\{\hat{\hat{f}}_{\theta}(X_{t})+R_{(i)}\}_{i=1}^{\lceil(n_{2}+1)(1-\alpha)\rceil},$ (5)
@@ 4 Experiments
The six ensemble analysis functions we consider include the pointwise average (EA), a weighted pointwise average (WA), the delta method (Delta), a pointwise linear model (LM), a Gaussian process regression using a neural-network Gaussian process (NNGP) kernel (GP) (Harris, Li and Sriver, 2023), and a deep convolutional neural network (CNN). WA and LM are the same models, except in WA, the weights are trained jointly, and there is no intercept term, whereas, in LM, the weights are learned independently at each spatial location. We used an NNGP kernel for the GP because it was shown to be more robust to covariate shift than standard isotropic kernels (Harris, Li and Sriver, 2023). For the CNN, we used five layers with a width of 64 and 3x3 convolutions with a skip connection between the 2nd and 4th layers. The CNN was trained with the Adam optimizer (Kingma and Ba, 2014) using a constant learning rate of $10^{-3}$ and no regularization, which was found to have no measurable impact on the results.
The three depth functions we consider include the $\ell_{\infty}$-Depth based on the $\ell_{\infty}$-norm, the Integrated Tukey depth (Tukey), and the (inverse) $\ell_{\infty}$ norm (norm), which is a non-depth based scoring rule. The $\ell_{\infty}$-Depth of a vector $r$ with respect to distribution $P$ is defined as $D_{\infty}(r,P)=1/(1+\mathbb{E}[||r-R||_{\infty}])$ where $R\sim P$. The Integrated Tukey depth averages the univariate Tukey depth $D_{\text{Tukey}}(r,P)=\mathbb{E}[\min\{\hat{F}_{x}(r(x),1-\hat{F}_{x}(r(x))\}]$ over each element of $r$. Here $\hat{F}_{x}$ denotes the empirical CDF of the residuals at location $x$. Finally, the $\ell_{\infty}$-norm simply computes $D(x)=1/(1+||x||_{\infty})$. We choose $\ell_{\infty}$-norm based depths and scoring rules, instead of the usual $\ell_{1}$ or $\ell_{2}$ norm based measures because $\ell_{\infty}$ was found to be more robust to changes in the distribution of the ensembles.
@@ 4.1 Coverage and sharpness
The white noise control run does not use any of the climate model data (Section 2.1), but instead simulates pure white noise fields from a standard Gaussian distribution. We sample “ensembles” of size $m=30$ on an $30\times 50$ grid, with 800 train observations, 200 validation observations, and 1000 test observations. We report the coverage on the test set averaged over the $m=30$ experiments. The white noise control runs are to verify that our approach works, regardless of the ensemble analysis and depth function, when exchangeability holds. After verifying on white noise, we then assess the empirical coverage on CMIP6 model data (TAS, TMAX, PR). TAS is the monthly average surface temperature, TMAX is the monthly maximum surface temperature, and PR is the (log) monthly total precipitation. TAS and PR are high-priority climate variables used extensively in climate change studies (IPCC), while TMAX allows us to assess our approach to extremal processes. Unlike the white noise runs, each of these variables exhibit distribution shift due to climate change (SSP2-4.5 forcings (O’Neill et al., 2016)).
@@
Figure 1 decomposes the results in Table 1 to show the coverage of each combination over time. We compute the coverage for each combination within (approximately) decadal blocks (2024-2030, 2030-2040,…,2090-2100) to illustrate how coverage is impacted by distribution shift in each variable. The $\ell_{\infty}$-Depth is highly stable around, or above, $0.9$ across all ensemble analysis functions and variable types. This suggests that is slightly conservative and thus relatively unaffected by distribution shift within each variable. As was seen in Table 1, the Tukey depth tends to be overly conservative and becomes more conservative over time in TAS and TMAX. The Infinity Norm is not robust against distribution shift, as seen by the dramatic loss of coverage in TAS and TMAX.
@@ 4.2 Uncertainty Quantification
Using our approach, we can construct a conformal prediction set and a conformal ensemble (Section 3.4). While the prediction set is theoretical, the conformal ensemble can be used in practice for uncertainty quantification (UQ). We compare the UQ skill of our approach, using each ensemble analysis and depth function combination, against the standard ensemble inter-model variability (IMV) and a quantile corrected version of the inter-model variability (IMV(BC)). The IMV and IMV(BC) are re-centered in each experiment from the ensemble mean to the ensemble analysis function’s projection.
@@
Table 2: Sliced Wasserstein distance between the projected distribution and the target process for each model analysis function and UQ method over the projection horizon (Apr. 2024 - Dec 2099). Each number represents the average SW distance across all perfect model experiments and the number in parenthesis is one standard error. Regardless of the analysis function, depth function, or climatic variable, conformal ensembles show lower projection error compared to the IMV and IMV(BC).
@@ 5.1 Historical validation
The second row of 5 shows the out-of-sample sliced Wasserstein distance between the projections (with uncertainty) and reanalysis fields on the held out data for PR. The first plot shows that using CE results uniformly lower SW numbers than either IMV or IMV(BC) across all ensemble analysis functions. Again, the Wasserstein diff maps are remarkably similar to each other and show that CE has a much better marginal approximation of the PR distribution in the Tropical Pacific near the Intertropical Convergence Zone (ITCZ). This region tends to feature variability on multiple time scales annual cycle (ITCZ), interseasonal variability (e.g., ENSO), and long-term trends due to global warming, all of which global climate models can struggle to emulate.
@@ 5.2 Forced response in the tails
Because each combination, GP + IMV, GP + IMV(BC), and GP + CE, produces the same mean prediction, we primarily investigate how each UQ method affects the upper (95%) and lower (5%) quantiles of the projected distribution. Specifically, we compute the 95% and 5% quantiles of each projection at each grid location over the 30 year projection period and compare these with the 95% and 5% quantiles of the historical reanalysis data at each grid location over the 30 year period Jan 1960 - Dec. 1989. Our goal is to show how each projection method forecasts changes in the distribution’s tails from a given reference period to show how each UQ method projects changes in the tails due to climate change.
@@
The top row of Figure 6 shows the 95% quantile change plots for TAS using the three projection methods. Each location on these plots shows the 95% quantile of the 30-year projection minus the 95% quantile of the 30-year historical period. Red areas indicate that the 95% quantile has increased over the reference period, while blue areas indicate that the 95% quantile has decreased. Both GP + IMV and GP + CE show heavy warming in the Arctic, around $6-8^{\circ}$ and $10-12^{\circ}$, respectively, and slight cooling in the Southern Ocean near Antarctica, around $0-2^{\circ}$. Both methods also show slight warming in the rest of the ocean. Overall, the changes in the 95% quantile between IMV and EA are broadly consistent with each other, with CE tending to show a more consistent increase with particularly extreme increases in the Arctic.
The bottom row of Figure 6 shows the 5% quantile change plots for TAS, in which GP + CE and GP + IMV heavily disagree. GP + CE shows consistent increases in the 5% quantile, indicating that it anticipates the lower range of the temperature distribution to increase dramatically over the next 30 years, particularly in high and low latitudes. This result is predicted from atmospheric physics based on reductions in the meridional temperature gradient, which climate models often struggle to capture (Lee, 2023). Together with the 95% quantile plot, this indicates that GP + CE is projecting an overall shift upwards in the temperature distribution. Conversely, GP + IMV is less certain about the projected distribution, with the 5% quantile maps showing sustained decreases, particularly in the Arctic ($2-4^{\circ}$).
@@ 6 Discussion and future work
Although our results showed that conformal ensembles have generally improved UQ skill compared to inter-model variability (even when bias-corrected), we view CE as a complementary approach to inter-model variability for climate uncertainty quantification. The IMV and a CE can be useful for assessing a multi-model ensemble’s unconditional and conditional uncertainty on a given projection task. Unconditional uncertainty is intrinsically useful because it allows us to quantify the models’ uncertainty. Comparing the two might be useful for investigating the predictability of different climatic variables (DelSole, 2004).
Furthermore, our conformal approach is intrinsically tied to the selected ensemble analysis function, depth function, calibration set, and even the reanalysis data product used. Our results showed that the choice of ensemble analysis function may not play a huge role, but more accurate functions generally result in tighter projection sets (Table 1). The choice of depth function does play a role in the robustness of the projection sets to violations of exchangeability, and proper depth functions (e.x. $\ell_{\infty}$-Depth or Tukey depth) should be preferred (Figure 1) over more typical scoring rules such as the $\ell_{\infty}$-Norm. Our sensitivity analysis showed that the calibration set can greatly affect UQ skill and that larger sets are not necessarily better (TAS and TMAX) due to the non-stationary nature of many climatic variables. Finally, if we swap ERA5 reanalysis data for a different reanalysis data product, such as NCEP (Kalnay et al., 1996) then our projections will likely change. In practice, tuning the calibration set size and considering multiple reanalysis data products may be necessary to marginalize these effects.

### v3

@@ 3.3 Conformal prediction sets for climate fields
Given a sufficiently adequate multi-model ensemble analysis function $f_{\theta}\in\mathcal{F}_{\Theta}$, where $f_{\theta}:\mathbb{R}^{q}\mapsto\mathbb{R}^{p}$, and a valid depth function $d:\mathbb{R}^{p}\times\mathcal{P}\mapsto[0,1]$ (Equation 4), we can implement the general ICP algorithm (Algorithm 1) for climate projection. Using the notation from Section 3.1, suppose we have a historical dataset $D_{\text{hist}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, an analysis function $f_{\theta}$, a depth function $d$, a confidence level $\alpha\in(0,1)$, and an out of sample input ensemble $X_{n+1}$. Algorithm 2 describes our procedure for computing $C_{\alpha}(X_{n+1})$ such that $P(Y_{n+1}\in C_{\alpha}(X_{n+1}))\geq 1-\alpha$.
@@
Algorithm 2 can be interpreted as “converting” the multi-model ensemble of climate model runs observed at time $n+1$, $X_{n+1}$, into an exact conformal prediction set $C_{\alpha}(X_{n+1})$. We take an inexact approximation of the projection distribution, model ensembles, and convert it into a constrained approximation, conformal ensembles. This approach is similar to Conformalized Quantile Regression (CQR) methods (Romano, Patterson and Candes, 2019; Kivaranovic, Johnson and Leeb, 2019) that convert arbitrary prediction quantiles into exact conformal quantiles. However, because the depth function $d(\cdot)$ measures outlyingness over the entire spatial domain simultaneously, the resulting projection sets $C_{\alpha}(X_{n+1})$ are jointly valid over $Y_{n+1}$ instead of only pointwise valid.
@@ 3.4 Conformal ensembles for uncertainty quantification
The standard approach to representing multivariate or functional prediction sets $C_{\alpha}(X_{n+1})$, generated by Algorithm 2, is to compute the pointwise minimum and maximum bounds that form a prediction band (Lei, Rinaldo and Wasserman, 2015; Diquigiovanni, Fontana and Vantini, 2021). Bands are commonly used because they are computationally efficient and form a convex hull around the level sets of $\mathcal{R}_{cal}$. However, because bands are a loose approximation, they contain many projections that do not belong to $C_{\alpha}(X_{n+1})$, and they discard information about the distribution of the ensemble. This makes bands less than ideal for climate projection, where we do not want our prediction sets to contain highly implausible climate fields and we want to assess the projection distribution.
@@
$E_{\alpha}(X_{t})=\{\hat{f}_{\theta}(X_{t})+R_{(i)}\}_{i=1}^{\lceil(n_{2}+1)(1-\alpha)\rceil},$ (5)
@@ 4 Experiments
The six ensemble analysis functions we consider include the pointwise average (EA), a weighted pointwise average (WA) (Bishop and Abramowitz, 2013), the delta method (Delta), a pointwise linear model (LM) (Abramowitz and Bishop, 2015), a Gaussian process regression using a neural-network Gaussian process (NNGP) kernel (GP) (Harris, Li and Sriver, 2023), and a deep convolutional neural network (CNN). WA and LM are the same models, except in WA, the weights are trained jointly, and there is no intercept term, whereas, in LM, the weights are learned independently at each spatial location. We used an NNGP kernel for the GP because it was shown to be more robust to covariate shift than standard isotropic kernels (Harris, Li and Sriver, 2023). For the CNN, we used five layers with a width of 64 and 3x3 convolutions with a skip connection between the 2nd and 4th layers. The CNN was trained with the Adam optimizer (Kingma and Ba, 2014) using a constant learning rate of $10^{-3}$ and no regularization, which was found to have no measurable impact on the results.
The three depth functions we consider include the $\ell_{\infty}$-Depth based on the $\ell_{\infty}$-norm, the Integrated Tukey depth (Tukey), and the (inverse) $\ell_{\infty}$ norm (norm), which is a non-depth based scoring rule. The $\ell_{\infty}$-Depth of a vector $r$ with respect to distribution $P$ is defined as $D_{\infty}(r,P)=1/(1+\mathbb{E}[||r-R||_{\infty}])$ where $R\sim P$. The Integrated Tukey depth averages the univariate Tukey depth $D_{\text{Tukey}}(r,P)=\mathbb{E}[\min\{\hat{F}_{x}(r(x),1-\hat{F}_{x}(r(x))\}]$ over each element of $r$. Here $\hat{F}_{x}$ denotes the empirical CDF of the residuals at location $x$. Finally, the $\ell_{\infty}$-norm simply computes $D(x)=1/(1+||x||_{\infty})$. We choose $\ell_{\infty}$-norm based depths and scoring rules, instead of the usual $\ell_{1}$ or $\ell_{2}$ norm based measures because $\ell_{\infty}$ was found to be more robust to changes in the distribution of the ensembles. In all cases, the resulting conformal ensemble could be generated in under a second on either an NVIDIA V100 or an NVIDIA 4700 GPU.
@@ 4.1 Coverage and sharpness
The white noise control run does not use any of the climate model data (Section 2.1), but instead simulates pure white noise fields from a standard Gaussian distribution. We sample “ensembles” of size $m=30$ on an 30×5030$\times$ 5030 × 50 grid, with 800 train observations, 200 validation observations, and 1000 test observations. We report the coverage on the test set averaged over the $m=30$ experiments. The white noise control runs are to verify that our approach works, regardless of the ensemble analysis and depth function, when exchangeability holds. After verifying on white noise, we then assess the empirical coverage on CMIP6 model data (TAS, TMAX, PR). TAS is the monthly average surface temperature, TMAX is the monthly maximum surface temperature, and PR is the (log) monthly total precipitation. TAS and PR are high-priority climate variables used extensively in climate change studies (IPCC), while TMAX allows us to assess our approach to extremal processes. Unlike the white noise runs, each of these variables exhibit distribution shift due to climate change (SSP2-4.5 forcings (O’Neill et al., 2016)).
@@
Figure 1 decomposes the results in Table 1 to show the coverage of each combination over time. We compute the coverage for each combination within (approximately) decadal blocks (2024-2030, 2030-2040,…,2090-2100) to illustrate how coverage is impacted by distribution shift in each variable. The $\ell_{\infty}$-Depth is highly stable around, or above, 0.90.90.9 across all ensemble analysis functions and variable types. This suggests that is slightly conservative and thus relatively unaffected by distribution shift within each variable. As was seen in Table 1, the Tukey depth tends to be overly conservative and becomes more conservative over time in TAS and TMAX. The Infinity Norm is not robust against distribution shift, as seen by the dramatic loss of coverage in TAS and TMAX.
@@ 4.2 Uncertainty Quantification
Using our approach, we can construct a conformal prediction set and a conformal ensemble (Section 3.4). While the prediction set is theoretical, the conformal ensemble can be used in practice for uncertainty quantification (UQ). We compare the UQ skill of our approach, using each ensemble analysis and depth function combination, against the standard ensemble inter-model variability (IMV) and a quantile corrected version of the inter-model variability (IMV(BC)) (Vrac and Friederichs, 2015). The IMV and IMV(BC) are re-centered in each experiment from the ensemble mean to the ensemble analysis function’s projection.
@@
Table 2: Sliced Wasserstein distance between the projected distribution and the target process for each model analysis function and UQ method over the projection horizon (Apr. 2024 - Dec 2099). Each number represents the average SW distance across all perfect model experiments and the number in parenthesis is one standard error. Regardless of the analysis function, depth function, or climatic variable, conformal ensembles (CE) show lower projection error compared to the IMV and IMV(BC).
@@ 4.3 Uncertainty quantification over time
Figure 2 also shows that CE is sensitive to distribution shift, as indicated by the steadily increasing sliced Wasserstein distance over time. However, because the rate of increase is the same across all UQ methods and analysis functions, this suggest that the source of this sensitivity is likely in the underlying analysis function. If the analysis function is sensitive to distribution shift then it will exhibit large errors which translates to poor distribution approximation, regardless of the chosen UQ method. I.e. a statistical model using CE based UQ does not show greater sensitivity to distribution shift than the baseline approach: an ensemble average using IMV based UQ.
@@ 5.1 Historical validation
The second row of 5 shows the out-of-sample sliced Wasserstein distance between the projections (with uncertainty) and reanalysis fields on the held out data for PR. The first plot shows that using CE results uniformly lower SW numbers than either IMV or IMV(BC) across all ensemble analysis functions. Again, the Wasserstein Diff. maps are remarkably similar to each other and show that CE has a much better marginal approximation of the PR distribution in the Tropical Pacific near the Intertropical Convergence Zone (ITCZ). This region tends to feature variability on multiple time scales annual cycle (ITCZ), interseasonal variability (e.g., ENSO), and long-term trends due to global warming, all of which global climate models can struggle to emulate.
@@ 5.2 Forced response in the tails
Because each combination, GP + IMV, GP + IMV(BC), and GP + CE, produces the same mean prediction, we primarily investigate how each UQ method affects the upper (95%) and lower (5%) quantiles of the projected distribution. Specifically, we compute the 95% and 5% quantiles of each projection at each grid location over the 30 year projection period and compare these with the 95% and 5% quantiles of the historical reanalysis data at each grid location over the 30 year period Jan 1960 - Dec. 1989. Our goal is to show how each projection method forecasts changes in the distribution’s tails, from a given reference period, to show how each UQ method projects changes in the tails due to climate change.
@@
The top row of Figure 6 shows the 95% quantile change plots for TAS using the three projection methods. Each location on these plots shows the 95% quantile of the 30-year projection minus the 95% quantile of the 30-year historical period. Red areas indicate that the 95% quantile has increased over the reference period, while blue areas indicate that the 95% quantile has decreased. Both GP + IMV and GP + CE show heavy warming in the Arctic, around $6-8^{\circ}$ and $10-12^{\circ}$, respectively, and slight cooling in the Southern Ocean near Antarctica, around $0-2^{\circ}$. Both methods also show slight warming in the rest of the ocean. Overall, the changes in the 95% quantile between GP + IMV and GP + CE are broadly consistent with each other, with GP + CE tending to show a more consistent increase with particularly extreme increases in the Arctic. GP + IMV(BC) shows similar spatial patterns of warming and cooling as GP + IMV but with a highly restricted range, projecting at most a $3^{\circ}$ warming. This effect is likely due to quantile mapping overfitting the tails of the historical data (Berg et al., 2024; Rahimi et al., 2024).
The bottom row of Figure 6 shows the 5% quantile change plots for TAS, in which GP + CE and GP + IMV heavily disagree. GP + CE shows consistent increases in the 5% quantile, indicating that it anticipates the lower range of the temperature distribution to increase dramatically over the next 30 years, particularly in high and low latitudes. This result is predicted from atmospheric physics based on reductions in the meridional temperature gradient, which climate models often struggle to capture (Lee, 2023). Together with the 95% quantile plot, this indicates that GP + CE is projecting an overall shift upwards in the temperature distribution. Conversely, GP + IMV is less certain about the projected distribution, with the 5% quantile maps showing sustained decreases, particularly in the Arctic ($2-4^{\circ}$). Again, GP + IMV(BC) is spatially consistent with GP + IMV but with a reduced range.
@@ 6 Discussion and future work
A major advantage of the conformal approach is that we do not require the data to be independent and identically distributed to rigorously guarantee the validity of the prediction sets. As long as the data are exchangeable, the bounds in Equation 1 will hold. This makes our method well suited to applications in climate simulations and reanalysis fields, which often exhibit strong spatiotemporal correlations. However, spatial and temporal correlations can still impact the empirical performance of our approach. We implicitly assumed that the data exhibit strong enough spatial structure to be treated as continuous surfaces so that functional data depth methods can be applied. Multi-model analysis functions that exploit spatial structure (E.g. CNNs and GPs) are likely to show higher accuracy and, therefore, more precise prediction sets. Temporal correlations could potentially impact the effective sample size of the calibration set, which could lead to excessive over-coverage (Equation 1), a phenomenon that we empirical observed (Table 1).
@@
Although our results showed that conformal ensembles have generally improved UQ skill compared to inter-model variability (even when bias-corrected), we view CE as a complementary approach to inter-model variability for climate uncertainty quantification. The IMV and a CE can be useful for assessing a multi-model ensemble’s unconditional and conditional uncertainty on a given projection task. Unconditional uncertainty is intrinsically useful because it allows us to quantify the models’ uncertainty. Comparing the two might be useful for investigating the predictability of different climatic variables (DelSole, 2004). CE could even be applied to single-model large initial conditional ensembles to investigate the internal variability of individual models. Focusing on a single model would allow us to isolate internal variability from other sources of variation such as inter-model variability.
Furthermore, our conformal approach is intrinsically tied to the selected ensemble analysis function, depth function, calibration set, and even the reanalysis data product used. Our results showed that the choice of ensemble analysis function may not play a huge role, but more accurate functions generally result in tighter projection sets (Table 1). The choice of depth function does play a role in the robustness of the projection sets to violations of exchangeability, and proper depth functions (e.x. $\ell_{\infty}$-Depth or Tukey depth) should be preferred (Figure 1) over more typical scoring rules such as the $\ell_{\infty}$-Norm. Our sensitivity analysis showed that the calibration set can greatly affect UQ skill and that larger sets are not necessarily better (TAS and TMAX) due to the non-stationary nature of many climatic variables. Finally, if we swap ERA5 reanalysis data for a different reanalysis data product, such as NCEP (Kalnay et al., 1996) then our projections will likely change. Exploring the potential for using multiple observational datasets, or higher frequency (daily) data, to constrain model ensembles in this framework remains an interesting direction for future work. Otherwise, in practice, tuning the calibration set size and considering multiple reanalysis data products may be necessary to marginalize these effects.